How do you solve the system equation?
6a+ 2b= 11
3a= 8b+ 1

Found 2 solutions by mananth, ewatrrr:
Answer by mananth(16949)

6a+ 2b= 11
3a= 8b+ 1
re write this equation.
3a-8b=1
multiply by -2 and it to first equation to eliminate x
-6a+16b=-2
add
6a-6a+2b+16b=11-2
18b=9
divide by 9
2b=1
/2
b = 1/2
..
plug value of b in second equation
3a=4+1
3a=5
divide by 3
a= 5/3
...
m.ananth@hotmail.ca

Answer by ewatrrr(24785)


Hi,
Adding the two equations in the system together & solving by eliminating one of the variables
6a+ 2b= 11
3a= 8b+ 1
arranging in like form
6a + 2b= 11
3a - 8b= 1
Multiplying first equation thru by 4
24a + 8b= 44
3a - 8b = 1
adding the two equation together (eliminating b variable)
27a = 45
a = 45/27 = 5/3
Solving for the b substituting a = 5/3
3a = 8b +1
5 = 8b + 1
4 = 8b
1/2 = b
checkiing the answer using one of the equation
24(5/3) + 8*(1/2) = 40 + 4 = 44